Campbell, a single taxpayer, earns $246,000 in taxable income and $3,600 in interest from an investment in State of New York bonds. (Use the U.S. tax rate schedule).

Required:

If Campbell earns an additional $17,000 of taxable income, what is her marginal tax rate on this income?

What is her marginal rate if, instead, she had $17,000 of additional deductions?
Answer #1

Taking the 2018 Federal Tax Rate structure,
(If Taxable Income is ranging from $200000 to $500000 - Tax Liability is $45689.50 + 35% in excess of $200000)

Campbell has Taxable Income of $246000 and will owe $61789.50 computed as follows :
Tax Liability = $45689.50 + 35% of ($246000 - $200000) = $61789.50

If Campbell earns an additional $17000 of taxable income, then the new tax liability will be $67739.50 computed as follows :
Taxable Income = $246000 + $17000 = $263000
Tax Liability = $45689.50 + 35% of ($263000 - $200000) = $67739.50

Marginal Tax Rate = Change in Tax / Change in Income
Marginal Tax Rate = ($67739.50 - $61789.50) / ($263000 - $246000)
Marginal Tax Rate = 35%

If Campbell had $17000 of additional tax deductions, then new tax liability will be $ computed as follows :
Taxable Income = $246000 - $17000 = $229000
Tax Liability = $45689.50 + 35% of ($229000 - $200000) = $55839.50

Marginal Tax Rate = ($55839.50 - $61789.50) / ($229000 - $246000)
Marginal Tax Rate = 35%

Conclusion :
Since in all cases the taxable income is in the same slab rate. Therefore, the marginal tax rate is same in both cases.